# Boosting Adverse Weather Crowd Counting via Multi-queue Contrastive Learning

### v2

@@ 4.3 JHU-Crowd++ Dataset
There are 4372 images and 1.51 million labels contained in the JHU-Crowd++ dataset [34]. Out of these, 2272 images were used for training, 500 images for validation, and the remaining 1600 images for testing. The advantage of JHU-Crowd++ is its inclusion of diverse scenes and environmental conditions, such as rain, snow and haze. It also provides weather condition labels for each image. Due to the rarity of adverse weather, the weather classes in JHU-Crowd++ are imbalanced. As is shown in Fig. 5 (a), the number of images under rain, snow, and haze conditions accounts for only 3%, 5%, and 4% of the total dataset, respectively.
Figure 5: The pie chart of the weather condition distribution of the JHU-Crowd++ dataset (a) and the GCC dataset (b).
@@
AWCC-Net [10] (ICCV 23) 58.0G+ 30M+
ConvNeXt [26] (CVPR 22) 27.0G 29M
MQCL (Ours) 31.2G 32M
Table 3 shows the performance of MQCL and other state-of-the-art methods on the JHU-Crowd++ dataset. Despite the challenges posed by the diverse scenes, complex and variable weather conditions and weather class imbalance in the JHU-Crowd++ dataset, MQCL achieves an improvement of 8.6% in MAE and 6.9% in RMSE under adverse weather conditions compared to the baseline ConvNeXt. This indicates that MQCL not only performs well under synthesized weather conditions but also effectively boosts the model’s robustness in real-world weather conditions. However, while MQCL achieved significant improvements over the baseline model ConvNeXt, its performance still slightly lags behind the current state-of-the-art algorithms. We summarize the reasons as follows: 1) as shown in Table 4, all methods outperforming MQCL, including CLTR, MAN, and AWCC-Net, have introduced Transformer-based modules, the computational complexity of which is significantly higher than MQCL. The MAE of AWCC-Net under adverse weather conditions is 9.5% lower than our method, but it incurs over 85.9% more FLOPs; 2) MQCL focuses on the design of learning strategies. It only introduces a lightweight refiner module without significantly altering the architecture of the backbone network. The visualizations of ConvNeXt and our MQCL are compared in Fig. 6.
Figure 6: Visualizations of our baseline ConvNeXt and our MQCL. The first row are the input images. The second and the third rows are the density maps predicted by ConvNeXt and our method, respectively. The images are sampled from the JHU dataset, including weather conditions of normal, haze, rain and snow.
4.4 GCC Dataset
GTA5 Crowd Counting (GCC) dataset [43] is a fully synthetic large scale dataset, the scenarios of which is derived from the computer game Grand Theft Auto V (GTA5). The dataset was initially proposed for pretraining the model and boosting the performance on the real-world datasets by finetuning or adaptation. Here, we use it as a benchmark to validate the performance of crowd counting models under adverse weather conditions on large-scale data. There are 15212 images and 7.63 million labels contained in the GCC dataset. Since each location in the dataset encompasses multiple perspectives, [43] provides three strategies for splitting the training and the testing set: random, cross-camera and cross-location. Here we choose the cross-location strategy, which means 75%/25% locations are utilized for training/testing, since this scheme is most effective in validating the robustness of the model against unknown scenes. The weather conditions in the GCC dataset are diverse, comprising a total of seven types: normal, cloudy, rainy, foggy, thunder, overcast and extra sunny. They are evidently unevenly distributed either and the pie chart of the weather distribution is shown in Fig. 5 (b). In our implementation, we randomly selected 10 locations out of the 75 in the training set as a validation set. And it is worth noting that we choose ConvNeXt-S rather than ConvNeXt-T as our backbone model due to the large-scale nature of the GCC dataset.
Table 5: Quantitative results comparing with the state-of-the-art methods on the GCC dataset. The numbers in parentheses represent the rankings of our method.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
CSRNet [19] (CVPR 18) 64.8 180.5 87.4 220.8
BL [28] (ICCV 19) 104.5 252.6 137.2 304.0
DM-Count [40] (NeurIPS 20) 78.4 207.6 103.5 248.5
KDMG [39] (PAMI 20) 94.6 225.0 124.0 280.5
MAN [22] (CVPR 22) 70.1 207.8 94.5 246.4
LDFNet [5] (TIP 24) 60.1 182.8 80.1 243.7
ConvNeXt-S [26] (CVPR 22) 65.3 184.5 85.5 221.7
MQCL (Ours) 59.7 (1) 175.5 (1) 79.7 (1) 210.5 (1)
The quantitative results of our method and several state-of-the-art methods are compared on the GCC dataset, as shown in Table 5. As can be seen, our MQCL achieves optimal performance under both normal and adverse weather conditions on all the metrics. Comparing to our baseline ConvNeXt, MQCL decreases the MAE/RMSE under normal and adverse weather conditions by 8.6%/4.9% and 6.8%/5.1%, respectively. This indicates that our MQCL can also outperforms other methods when trained on relatively large-scale datasets and more importantly, our multi-queue MoCo strategy performs well even in the case of higher number of sub queues (7 for example).
Table 6: The significance of contrastive learning in our method. “Refine w/o CL” adopts a similar approach to MQCL, except that $\mathcal{L}_{contra1}$ and $\mathcal{L}_{contra2}$ are removed from the loss functions, respectively.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Encoder and decoder only 52.7 154.9 105.1 561.4
Refine w/o CL 56.2 163.1 107.1 585.6
MQCL (Ours) 50.5 152.0 96.5 522.8
4.5 Key Issues and Discussions
Strategy of positive/negative selection. As mentioned above, there are two optional strategies in the WRL stage: 1) treating samples with the same weather label as positive examples; 2) treating samples originating from the same image as positive examples; We conduct experiments on the JHU-Crowd++ dataset to compare these two strategies. The model performance after the WRL stage using these two strategies is shown in Table LABEL:tbl:strategy_comparison. Strategy 2 outperforms Strategy 1 in both normal and adverse conditions. We attribute this to the following reasons: 1) even if two samples share the same weather label, their weather conditions may still vary significantly. The practice of minimizing all the representations with the same label is not in line with the target of us; 2) samples from the same image not only share the same weather conditions but also the same scene. Strategy 2 has the potential to enhance the model’s scene recognition capabilities.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Memory Bank 53.7 157.6 104.9 531.1
Single-queue MoCo 54.6 166.5 107.9 589.7
Multi-queue MoCo (Ours) 50.5 152.0 96.5 522.8
The storage strategy for vectors K. There are three optional storage strategies for vectors K: 1) memory bank [44]; 2) single-queue MoCo [8]; 3) multi-queue MoCo (ours). Extensive experiments are conducted on the JHU-Crowd++ dataset to compare the above strategies. The t-SNE [37] visualization of the vectors Q after the WRL stage on the JHU-Crowd++ dataset is shown in Fig. 7. The memory bank strategy suffers from a severe lack of discriminative capacity. The representations learned by the single-queue strategy are also not discriminative enough, especially near the rain weather representations in Fig. 7(b). The multi-queue strategy does not suffer from the aforementioned issues. The significantly dispersed clusters shown Fig. 7(c) illustrate that the properties learned from the multi-queue MoCo represent weather information well. The performance comparison shown in Table 7 also demonstrates that the proposed multi-queue MoCo can effectively address the class imbalance problem. We attribute the phenomenon to the following reasons: 1) while the memory bank can store a large number of samples with minimal memory consumption, it does not employ a stable strategy to update the encoder, and the sample update frequency is too low, resulting in poor sample consistency; 2) although single-queue MoCo addresses the issue of poor sample consistency by introducing a queue and momentum update strategy, the class imbalance problem in the dataset leads to inconsistencies between the data distribution in the queue and the uniform distribution prior; 3) multi-queue MoCo not only retains the advantages of high sample consistency but also greatly alleviates the problem of inconsistency between data distribution and the uniform prior. Thus, it achieves the best performance.
Table 8: Ablation study on the JHU-Crowd++ dataset.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Encoder and decoder only 52.7 154.9 105.1 561.4
+ Multi-queue MoCo 51.2 152.6 97.2 524.6
+ Refiner 50.5 152.0 96.5 522.8
Ablation study. Ablation studies are performed on the JHU-Crowd++ dataset and the quantitative results are shown in Table 8. We start with the baseline of the end-to-end model, i.e., only the encoder and decoder. First, the effectiveness of multi-queue MoCo is tested. An improvement of 2.8%/1.5% and 7.5%/6.6% in MAE/RMSE under normal and adverse weather is achieved compared to the baseline. From this, it can be seen that most of the performance improvements under adverse weather conditions come from the representation learning in the WRL stage. This verifies the effectiveness of the proposed multi-queue MoCo for enhancing robustness under adverse weather conditions. Additionally, the performance improvement under normal weather conditions corroborates the earlier analysis that the strategy that treats different augmentations from the same image as positive samples can aid in strengthening the scene recognition capabilities of the model. Subsequently, the refiner is added, and the best performance is achieved, with a reduction of 1.4%/0.4% in MAE/RMSE under normal weather conditions and 0.7%/0.3% under adverse weather conditions, respectively. This demonstrates that the refiner is capable of converting adverse weather representations to the normal domain, enabling the decoder to focus on a single domain.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
KCL [13] (ICLR 20) 51.1 154.3 97.9 564.6
BCL [46] (CVPR 22) 50.5 154.3 97.2 524.2
TSC [18] (CVPR 22) 50.8 152.1 96.4 522.4
SBCL [9] (ICCV 23) 54.0 159.2 101.1 563.0
Multi-queue MoCo (Ours) 50.5 152.0 96.5 522.8
@@ 4.5 Key Issues and Discussions
Table 10: The performance of MQCL on the JHU-Crowd++ dataset under different encoder configurations.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
VGG [33] 68.2 221.4 115.3 632.0
VGG + MQCL 67.2 209.2 101.4 533.4
CSRNet [19] 64.3 223.5 114.8 603.9
CSRNet + MQCL 63.1 207.7 99.8 540.2
Swin [25] 57.8 171.8 103.5 571.7
Swin + MQCL 54.6 163.8 99.5 558.8
ConvNeXt [26] 52.7 154.9 105.1 561.4
MQCL 50.5 152.0 96.5 522.8
Performance under different encoder configurations. Since the focus of MQCL is the designing of learning strategies, it is necessary to validate the effectiveness of MQCL under different network architectures. We conduct experiments on the JHU-Crowd++ dataset, where the encoder of MQCL is replaced with other models, including simple CNN models like VGG [33] and CSRNet [19] and a Transformer-based model Swin-T [25]. The result is shown in Table 10. The structure of the base models (ConvNeXt, CSRNet, VGG, Swin) keep consistent of the original model, only the classification head is replaced with the same decoder used in MQCL. Clearly, applying any encoder structure to MQCL results in significantly better performance compared to the corresponding baseline. This demonstrates the versatility of the learning methods in MQCL.
Table 11: Generalization ability of ConvNeXt-T trained on the NWPU-Normal dataset and the NWPU-Weather dataset on the test set of the JHU-Crowd++ dataset.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
NWPU-Normal to JHU-Crowd 65.9 184.4 144.7 705.2
NWPU-Weather to JHU-Crowd 63.9 163.6 142.8 497.2
The significance of the NWPU-Weather dataset. We think the significance of the NWPU-Weather dataset is two-fold. Firstly, since there is currently only one fully real-world adverse weather crowd counting dataset, and the GCC dataset is not based on real-world scenarios. We think that these two datasets are insufficient to validate the model’s performance under adverse weather conditions. Therefore, we synthesize the NWPU-Weather dataset based on real-world scenarios as a supplement. Additionally, beyond this, there is a question: how is the generalization ability of the model trained on the synthetic NWPU-Weather dataset? If the model trained on the NWPU-Weather dataset can’t be generalized to the real world, the contribution of the synthetic dataset is limited. Thus we conduct experiments to validate this question.
In the experiment setup, we prepare two datasets as source-domain datasets. One is our proposed NWPU-Weather dataset. And another, for comparison, we replace all the adverse-weather images in the NWPU-Weather dataset with its corresponding normal-weather raw images. Here we call it NWPU-Normal. To test the generalization ability of the models, the target domain is set as the JHU-Crowd dataset since it is fully real-world. We train ConvNeXt-T on the training set of the two source-domain datasets respectively and directly test their performance on the testing set of the target domain dataset without any finetuning.
As is shown in Fig. 11, compared to the model trained on the NWPU-Normal dataset, that trained on our synthesised NWPU-Weather dataset is 3.0%/11.3% and 1.3%/29.5% stronger in MAE/RMSE under normal and adverse weather, respectively. This validates the significance of our NWPU-Weather dataset: more diversified weather conditions in the training data imparts better robustness to the model.

### v3

@@ 4.2 JHU-Crowd++ Dataset
There are 4372 images and 1.51 million labels contained in the JHU-Crowd++ dataset [30]. Out of these, 2272 images were used for training, 500 images for validation, and the remaining 1600 images for testing. The advantage of JHU-Crowd++ is its inclusion of diverse scenes and environmental conditions, such as rain, snow and haze. It also provides weather condition labels for each image. Due to the rarity of adverse weather, the weather classes in JHU-Crowd++ are imbalanced. As is shown in Figure 1 (a), the number of images under rain, snow, and haze conditions accounts for only 3%, 5%, and 4% of the total dataset, respectively.
Table 1 shows the performance of MQCL and other state-of-the-art methods on the JHU-Crowd++ dataset. VGG-16, Swin-T and ConvNeXt-T in the table stand for the baseline models, which have the same structure with their original image classification models, except for replacing the classification heads with crowd counting heads. Despite the challenges posed by the diverse scenes, complex and variable weather conditions and weather class imbalance in the JHU-Crowd++ dataset, MQCL achieves significant performance improvements regardless of the backbone used. For example, MQCL achieves an improvement of 21.8% in MAE and 22.8% in RMSE under adverse weather conditions compared to the baseline ConvNeXt-T. Moreover, MQCL can also improve the performance under normal conditions, decreasing MAE and RMSE by 8.7% and 3.2% compared to baseline ConvNeXt-T, respectively. We believe that this can be attributed to the contrastive learning strategy employed in this paper, which considers augmentations from the same image as positive examples. This strategy not only assists the model in weather perception but also strengthens the model’s ability to recognize different scenes. Sampling from the same image ensures that the positive examples not only share the same weather conditions but also possess similar scene characteristics. On the other hand, owing to the superior performance in adverse weather conditions, MQCL surpasses state-of-the-art models such as MAN and AWCC-Net in overall test-set performance.
In addition, we conducted experimental verification of two simple methods mentioned in the Introduction, both based on the ConvNeXt-T baseline. The approach of training a separate model for each weather condition corresponds to “MultiModel” in the table. For this method, ResNet-50 [7] is used as the weather classifier. It is evident that the limited number of adverse weather images in the training set is insufficient to support the training of high-performance models, leading to worse performance under adverse weather compared to the baseline. Moreover, occasional misclassifications by the weather classifier also negatively impact performance under normal weather conditions. The approach of performing image restoration before counting corresponds to “Restore” in the table, where we adopt the Unified model [4] for restoration. Under this method, the performance under adverse weather conditions shows some improvement, but the gains are significantly smaller compared to those achieved by MQCL. The visualizations of ConvNeXt-T and our MQCL are compared in Figure 4.
Figure 4: Visualizations of our baseline ConvNeXt-T and our MQCL. The first row are the input images. The second and the third rows are the density maps predicted by ConvNeXt and our method, respectively. The images are sampled from the JHU dataset, including weather conditions of normal, haze, rain and snow.
4.3 GCC Dataset
GTA5 Crowd Counting (GCC) dataset [37] is a fully synthetic large scale dataset, the scenarios of which is derived from the computer game Grand Theft Auto V (GTA5). The dataset was initially proposed for pretraining the model and boosting the performance on the real-world datasets by finetuning or adaptation. Here, we use it as a benchmark to validate the performance of crowd counting models under adverse weather conditions on large-scale data. There are 15212 images and 7.63 million labels contained in the GCC dataset. Since each location in the dataset encompasses multiple cameras, [37] provides three strategies for splitting the training and the testing set: random, cross-camera and cross-location. Here we choose the cross-location strategy, which means 75%/25% locations are utilized for training/testing, since this scheme is most effective in validating the robustness of the model against unknown scenes. The weather conditions in the GCC dataset are diverse, comprising a total of seven types: normal, cloudy, rainy, foggy, thunder, overcast and extra sunny. They are evidently unevenly distributed either and the pie chart of the weather distribution is shown in Figure 1 (b). In our implementation, we randomly selected 10 locations out of the 75 in the training set as a validation set.
Table 2: Quantitative results comparing with the state-of-the-art methods on the GCC dataset. The best performance is shown in bold and the second best is shown in underlined.
Method Normal Adverse Total
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
CSRNet [17] (CVPR 18) 64.8 180.5 87.4 220.8 75.5 200.7
BL [25] (ICCV 19) 104.5 252.6 137.2 304.0 120.0 278.2
DM-Count [35] (NeurIPS 20) 78.4 207.6 103.5 248.5 90.3 227.9
KDMG [33] (PAMI 20) 94.6 225.0 124.0 280.5 108.6 252.9
MAN [19] (CVPR 22) 70.1 207.8 94.5 246.4 81.7 227.0
LDFNet [5] (TIP 24) 60.1 182.8 80.1 243.7 69.6 213.9
APGCC [2] (ECCV 24) 63.2 175.9 87.2 258.5 74.6 219.1
ConvNeXt-T [23] (CVPR 22) 65.3 184.5 85.5 221.7 74.9 203.0
MQCL + ConvNeXt-T 59.7 175.5 79.7 210.5 69.2 192.9
We conducted quantitative experiments on the GCC dataset using MQCL with a ConvNeXt-T backbone. Note that fewer comparison methods is included on this dataset because all performance results are obtained from our own training, as shown in Table 2. Our MQCL achieves optimal performance under both normal and adverse weather conditions on all the metrics. Compared to our baseline ConvNeXt-T, MQCL decreases the MAE/RMSE under normal and adverse weather conditions by 8.6%/4.9% and 6.8%/5.1%, respectively. This indicates that our MQCL can also outperforms other methods when trained on relatively large-scale datasets and more importantly, our multi-queue MoCo strategy performs well even in the case of higher number of sub queues (7 for example).
4.4 Key Issues and Discussions
Table 3: Performance comparison of the two optional strategies on the JHU-Crowd++ dataset. Strategy 1 treats samples with the same weather label as positive examples and strategy 2 treats samples originating from the same image as positive examples.
Method Normal Adverse Total
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Strategy 1 51.4 153.0 89.7 470.4 56.0 216.9
Strategy 2 (Ours) 48.1 150.0 82.2 433.5 52.1 205.5
Strategy of positive/negative selection. As mentioned above, there are two optional strategies in the WRL stage: 1) treating samples with the same weather label as positive examples; 2) treating samples originating from the same image as positive examples; We conduct experiments on the JHU-Crowd++ dataset to compare these two strategies. The model performance under these two strategies is shown in Table 3. Strategy 2 outperforms Strategy 1 in both normal and adverse conditions. We attribute this to the following reasons: 1) even if two samples share the same weather label, their weather conditions may still vary significantly. The practice of minimizing all the representations with the same label is not in line with the target of us; 2) samples from the same image not only share the same weather conditions but also the same scene. Strategy 2 has the potential to enhance the model’s scene recognition capabilities.
Method Normal Adverse Total
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Memory Bank 53.7 157.6 104.9 531.1 59.8 235.7
Single-queue MoCo 53.6 166.5 98.9 452.7 59.0 221.1
Multi-queue MoCo (Ours) 48.1 150.0 82.2 433.5 52.1 205.5
The storage strategy for vectors K. There are three optional storage strategies for vectors K: 1) memory bank [38]; 2) single-queue MoCo [8]; 3) multi-queue MoCo (ours). Extensive experiments are conducted on the JHU-Crowd++ dataset to compare the above strategies. The t-SNE [32] visualization of the vectors Q after the WRL stage on the JHU-Crowd++ dataset is shown in Figure 5. The memory bank strategy suffers from a severe lack of discriminative capacity. The representations learned by the single-queue strategy are also not discriminative enough, especially near the rain weather representations in Figure 5(b). The multi-queue strategy does not suffer from the aforementioned issues. The significantly dispersed clusters shown Figure 5(c) illustrate that the properties learned from the multi-queue MoCo represent weather information well. The performance comparison shown in Table 4 also demonstrates that the proposed multi-queue MoCo can effectively address the class imbalance problem. We attribute the phenomenon to the following reasons: 1) while the memory bank can store a large number of samples with minimal memory consumption, it does not employ a stable strategy to update the encoder, and the sample update frequency is too low, resulting in poor sample consistency; 2) although single-queue MoCo addresses the issue of poor sample consistency by introducing a queue and momentum update strategy, the class imbalance problem in the dataset leads to inconsistencies between the data distribution in the queue and the uniform distribution prior; 3) multi-queue MoCo not only retains the advantages of high sample consistency but also greatly alleviates the problem of inconsistency between data distribution and the uniform prior. Thus, it achieves the best performance.
Table 5: Ablation study on the JHU-Crowd++ dataset.
WRL CRR Normal Adverse Total
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
52.7 154.9 105.1 561.4 59.0 242.4
49.1 152.6 87.2 458.5 53.6 213.5
48.1 150.0 82.2 433.5 52.1 205.5
Ablation study. Ablation studies are performed on the JHU-Crowd++ dataset and the quantitative results are shown in Table 5. We start with the baseline of ConvNeXt-T. First, the effectiveness of the WRL stage is tested. An improvement of 6.8%/1.5% and 17.0%/18.3% in MAE/RMSE under normal and adverse weather is achieved compared to the baseline. From this, it can be seen that most of the performance improvements under adverse weather conditions come from the representation learning in the WRL stage. This verifies the effectiveness of the proposed multi-queue MoCo for enhancing robustness under adverse weather conditions. Additionally, the performance improvement under normal weather conditions corroborates the earlier analysis that the strategy that treats different augmentations from the same image as positive samples can aid in strengthening the scene recognition capabilities of the model. Subsequently, the CRR stage is added, and the best performance is achieved, with a reduction of 2.0%/1.7% in MAE/RMSE under normal weather conditions and 5.7%/5.5% under adverse weather conditions, respectively. This demonstrates that the refiner is capable of converting adverse weather representations to the normal domain, enabling the decoder to focus on a single domain.
Method Normal Adverse Total
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
KCL [12] (ICLR 20) 48.7 154.3 83.9 464.6 52.9 216.1
BCL [40] (CVPR 22) 48.1 154.3 83.2 524.2 52.3 231.9
TSC [16] (CVPR 22) 48.8 152.1 82.0 437.4 52.8 207.9
SBCL [9] (ICCV 23) 52.0 159.2 87.1 473.0 56.2 221.4
Multi-queue MoCo (Ours) 48.1 150.0 82.2 433.5 52.1 205.5
@@ 4.4 Key Issues and Discussions
AWCC-Net [10] (ICCV 23) 58.0G 30M
Gramformer [20] (AAAI 24) 60.9G 29M
ConvNeXt-T [23] (CVPR 22) 27.0G 29M
MQCL + ConvNeXt-T 31.2G 32M
Computation complexity and number of parameters. As shown in Table 7, compared to the baseline, MQCL introduces only a light-weight refiner, resulting in only 4.2GFLOPs in computational complexity and 3M parameters. Compared to state-of-the-art methods, MQCL has similar parameter counts to Gramformer and AWCC-Net, while requiring only half their computational cost. CLTR has higher requirements on both computational and memory cost than MQCL. Superior performance to state-of-the-art methods with lower computational costs demonstrates MQCL’s successful achievement of light-weight design objective.